# The Winner of the NFL Draft is Not Necessarily Cursed

### v1

@@ Abstract
Football analysts traditionally determine the relative value of draft picks by average future player value at each draft position. One implication is the loser’s curse: top draft picks belonging to last year’s worst teams produce less surplus value on average than draft picks later in the first round belonging to better teams. Additionally, these valuations do not match the valuation implied by the trade market. Either general managers are making terrible trades on average, or there is a sound economic reason for the discrepancy; we are partial to the latter explanation. Traditional analyses don’t consider that variance in performance decays convexly accross the draft, causing eliteness (e.g., right tail probability) to decay much more steeply than expected value. Because elite players have an outsize influence on winning the Super Bowl, we suspect general managers value performance nonlinearly, placing exponentially higher value on players as their eliteness increases. Draft curves that account for this closely resemble the trade market. Additionally, we create draft curves that adjust for position via a novel Bayesian hierarchical Beta regression model. We find that if you are interested in an elite quarterback, there is no loser’s curse.
Keywords: Decision making under uncertainty, applications and case studies, statistics in sports, NFL Draft, Beta regression, Bayesian hierarchical model
@@ 1 Introduction
1.1 Background information
The NFL Draft. Players joining the National Football League (NFL) are allocated to teams via an annual draft. Teams take turns selecting players in an order determined by their win-loss records from the previous year. In each of seven rounds, teams select players in increasing order of the previous year’s win-loss record. Drafting begins with last year’s worst team and ends with last year’s champion (with some minor exceptions). Each drafted player is then signed to a four-year first contract, also known as the rookie contract. Teams can choose to add a fifth year to the contracts of first-round selections, known as a fifth-year team option.
Player quality decays convexly over the course of the draft. Players drafted in the first round often turn out to be solid or excellent NFL players. Players drafted in the seventh round rarely turn out to be solid NFL players. Reflecting this trend, a drafted player’s first contract salary is determined by his draft position. First contract compensation decreases convexly over the first few rounds of the draft, leveling off after the fourth round. First contract compensation is essentially deterministic, although drafted rookies sometimes negotiate elements of the contract such as performance bonuses or injury guarantees. These rookie contracts count towards the NFL salary cap, a rule that determines the maximum amount any team can pay its players in a given year. Over time the cap has increased from \$123 million in 2013 to \$224.8 million in 2023. But, rookies are generally paid much less than veteran players who sign in free agency. In free agency, NFL players who are not bound by a contract can sign a new contract with any team.
The Market for NFL Draft Picks. Trading draft picks is a big part of football. NFL teams often find themselves wanting to trade draft picks, and such trades typically consist of current picks, future picks, and/or active players. A general manager may want to trade up to draft a particular player before he is taken by another team. For example, in 2017 the Kansas City Chiefs sent a first-round pick (the $27^{th}$ pick), a third-round pick (the $91^{st}$ pick), and its 2018 first-round pick to the Buffalo Bills in exchange for the $10^{th}$ pick, which it used just minutes after the trade to draft quarterback Patrick Mahomes. A general manager may also want to trade up to be able to draft players of a certain caliber at a particular position. For example, in 2023 the Carolina Panthers sent a first-round pick (the $9^{th}$ pick), a second-round pick (the $61^{st}$ pick), its 2024 first-round pick, its 2025 second-round pick, and wide receiver DJ Moore to the Chicago Bears for the $1^{st}$ overall pick, with the intention of drafting an elite quarterback.
In 2017, how did the Chiefs and the Bills decide that the draft-day trade was worth executing? More generally, if a team wants to trade for a particular draft pick, which picks should it offer in exchange? And if a team is offered a bundle of draft picks in exchange for another bundle of draft picks, should it accept or reject the trade? We are interested in the relative value of draft picks.
1.2 Traditional draft value curves
Football analysts who calculate the relative value of draft picks typically display their valuations as charts or curves, plotting value against draft position (e.g., see Figure 1 later in this section).
Jimmy Johnson’s Chart. In the 1990s, at the request of Dallas Cowboys coach Jimmy Johnson, vice president Mike McCoy created the first draft value curve based on gut instinct and looking at past trades. Known today as the Jimmy Johnson chart,111 https://www.the33rdteam.com/jimmy-johnsons-draft-trade-value-chart-set-standard-but-clearly-outdated/ it assigns a numerical point value to each draft position $1,...,256$. The first pick is worth the most, and the curve decays steeply and convexly. Jimmy Johnson’s formulation, and draft curves today, are additive: the value of a bundle of picks is the sum of the values of each individual pick. A trade is considered fair if the sum of the values of the two sides of the trade is equal. For example, consider the 2017 draft trade between the Chiefs and Bills, discussed earlier. According to Jimmy Johnson’s chart, the Bills gained value by accepting the trade: the Chiefs three picks are worth about 1,816 points,222 The Chiefs three picks are worth 680, 136, and 1000 points, respectively, supposing the Chiefs’ 2018 first-round pick is in the middle of the first-round at pick 16, which is reasonable considering the Chiefs made the playoffs 3 of the 4 years before 2017. more than the 1,300 point value of the Bills’ $10^{th}$ pick.
Though Jimmy Johnson’s chart is ad hoc and nonrigorous, general managers used it for decades, and some teams still use it today. More recently, Massey and Thaler, (2013) created rigorous data-driven draft value curves. In particular, they considered the expected value of a drafted player’s performance, and performance relative to cost, as a function of draft position.
@@ 1.2 Traditional draft value curves
Measuring player performance. Defining the value of a player’s performance in the first place takes care and consideration. Each play in American football consists of a complex interplay between 22 players of varying roles interacting in continuous time and space. This makes it difficult to measure how a player’s individual performance impacts the ultimate outcome of a play. For instance, a linebacker may impact a play simply by occupying space. He could cause the ball carrier to move in a different direction, and there is no easy way to measure the impact of that action. There is no easy or industry-standard way of measuring player performance on a given play in American football, particularly in a way that measures players of all positions on the same scale. Hence, analysts use performance proxies. Each measure is imperfect and has its own drawbacks. We describe many of these performance measures in detail in Appendix B. Throughout this paper, for simplicity, we use one performance outcome variable, detailed below.
@@
In free agency, NFL players who are not bound by a contract can sign a new contract with any team. We believe the free agent market is relatively efficient and thus reflects the performance value of players. Therefore, following Baldwin, (2024), in this study we measure performance value by the average salary per year of a contract signed in free agency. We consider salary as a percentage of the cap, known as APY cap percentage, to adjust for inflation.
@@
Football analysts, beginning with Massey and Thaler, (2013), have been interested in particular in a player’s first contract performance value rather than his career performance value. This is because an NFL player is paid particularly cheaply during his first contract. A rookie’s first contract is determined by his draft position and can be much lower than the value his performance provides, creating a first contract “surplus value.” Teams can take advantage of this extra cap space by signing good veteran players in free agency. The drawback of focusing solely on a player’s first contract is that the very best players stay with the same team across multiple contracts, providing value across their entire careers (e.g., Tom Brady). Continuing with the precedent set by Massey and Thaler, (2013) and subsequent analyses, in this study we consider just first contract performance and surplus value.
@@
Thus, throughout this paper, we measure a player’s first contract performance by his second contract’s average salary per year as a percentage of the cap (APY cap percentage). This is a reflection of a player’s first contract performance value since he signs his second contract in free agency just after his first contract ends. Players who don’t sign a second contract have 0 performance value. We acknolwedge that free agent APY cap percentage is an imperfect measure of a player’s performance (see Appendix B) and first contract value is an incomplete measure of the total value he provides to his team (as it excludes the rest of his career). Nonetheless, the focus of this paper is to value draft picks given a performance measure, not to choose the best performance measure. We leave further analysis of various measures of performance to future work.
@@
Traditional expected value curves. Now, we discuss the methodology underlying traditional data-driven draft curves. Formally, prior to the draft, we do not know how well a player drafted at pick $x\in\{1,...,256\}$ will perform in the NFL. Thus, we think of the performance outcome $Y$ associated with pick $x$ as a random variable, denoted by a capital letter. Recall that, throughout this paper, we let the realization of $Y$ be a player’s annualized second contract value relative to the salary cap (APY cap percentage). Then, from the dataset $\{(x_{i},Y_{i})\}$ of all recent draft picks in NFL history, we want to learn the de-noised relationship between pick number $x$ and performance outcome $Y$. Traditional analyses use a standard approach but with different measures of performance, detailed in Appendix B (Massey and Thaler,, 2013; Stuart,, 2012; Fitzgerald and Spielberger,, 2024; Pro Football Focus,, 2024; Baldwin,, 2024). The de-noised relationship they estimate from data is the expected value of $Y$ given $x$.
@@
This traditional approach proceeds as follows. First, fit the expected performance curve, or the conditional mean $x\mapsto{\mathbb{E}}[Y|x]$. Second, calculate the cost curve, $x\mapsto{\mathrm{cost}}(x)$. Cost is first contract compensation, which is essentially a deterministic function of $x$. Third, calculate the expected surplus curve, or the expected value of the difference between the performance and cost, $x\mapsto{\mathbb{E}}[Y-{\mathrm{cost}}(x)|x]$. Finally, normalize each of these curves so that the value of the first pick is 1. Analysts typically normalize by dividing by the value of the first pick.
@@
We fit these traditional draft value curves using our outcome variable $Y$ (second contract APY cap percentage), which we visualize in Figure 1. We use spline regression to fit the expected performance curve from a dataset of all draft picks from 2013 to 2023 (Ho and Carl,, 2024). We calculate the cost curve as in Baldwin, (2024): beginning with the actual 2023 salary cap (about \$$\$225$ million) and assuming a cap growth rate of 7% over the next three seasons (2024, 2025, and 2026), we convert the total dollar amount of the rookie contract into a percentage of the cap. Then, for each draft pick we average the compensation values across the four seasons. We calculate the expected surplus curve as the difference between the expected performance curve and cost curve.
@@
Figure 1: Each curve is a graph of the value $v(x)$ of draft position $x$ relative to the first pick. In blue is the expected performance curve, proportional to $x\mapsto{\mathbb{E}}[Y|x]$. In red is the cost curve, proportional to $x\mapsto{\mathrm{cost}}(x)$. In green in the expected surplus curve, proportional to $x\mapsto{\mathbb{E}}[Y-{\mathrm{cost}}(x)|x]$. In orange is the draft value curve implied by the trade market. In pink is Jimmy Johnson’s curve.
@@
Expected performance value (blue) is less steep than cost (red), indicating that players are underpaid on average in their first contracts. This produces positive expected surplus value (green) across most of the draft. Expected surplus value (green) is larger than 1 throughout the first round, peaking in the middle of the first round. This surplus spike is the “loser’s curse” from Massey and Thaler, (2013): top draft picks, belonging to last year’s worst teams, produce less surplus value on average than draft picks later in the first round that belong to better teams.
We also visualize Jimmy Johnson’s draft curve (pink) and the draft curve fit from the trade market (orange). Devised in Massey and Thaler, (2013), the fitted trade market curve is a singular draft value curve that best fits the dataset of all observed trades. We derive this curve in detail in Appendix C.

### v2

# The Loser’s Curse Revisited: A Case Study of the Critical Role of Specifying a Utility Function

@@ Abstract
A longstanding question in the judgment and decision making literature is whether experts, even in high-stakes environments, exhibit the same cognitive biases observed in controlled experiments with inexperienced participants. In their seminal work, Massey and Thaler (2013) provide a notable example of bias and irrationality in expert decision making: general managers’ behavior in the National Football League draft pick trade market. They argue that general managers systematically overvalue top draft picks, which generate less surplus value on average than later first-round picks, a phenomenon known as the loser’s curse. Their conclusion hinges on the assumption that general managers should use expected surplus value as their utility function for evaluating draft picks. This assumption, however, is neither explicitly justified nor necessarily aligned with the strategic complexities of constructing a National Football League roster. In this paper, we challenge their framework by considering alternative utility functions, particularly those that emphasize the acquisition of transformational players––those capable of dramatically increasing a team’s chances of winning the Super Bowl. Under a decision rule that prioritizes the probability of acquiring elite players, which we construct from a novel Bayesian hierarchical Beta regression model, general managers’ draft trade behavior appears rational rather than systematically flawed. More broadly, our findings highlight the critical role of carefully specifying a utility function when evaluating the quality of decisions.
Keywords: Decision making under uncertainty, applications and case studies, statistics in sports, NFL Draft, Beta regression, Bayesian hierarchical modeling
@@ 1 Introduction
A longstanding question in the judgment and decision making literature is whether experts, when making high-stakes decisions, exhibit the same cognitive biases and irrational behaviors documented in studies involving inexperienced participants in low-stakes environments. Critics of this literature argue that biases may be overstated, as most studies do not analyze decision making in real-world, high-stakes settings. Massey and Thaler, (2013) provide a notable exception, applying behavioral economic principles to a context––the National Football League (NFL) draft––where decision makers, namely NFL general managers, are highly experienced, financially incentivized, and have access to extensive data and learning opportunities.
Massey and Thaler’s study has had a significant impact not just in sports analytics but in the broader decision making literature, reinforcing the notion that systematic biases can persist even in expert settings. They argue that general managers systematically overpay for top draft picks despite evidence that these selections generate less “surplus value”––the difference between the value of a player’s performance and his cost––than later first-round picks, on average. This so-called “loser’s curse,” widely cited as evidence of irrational decision making, is the subject of their ultimate striking conclusion: “top draft picks are significantly overvalued in a manner that is inconsistent with rational expectations and efficient markets, and consistent with psychological research” (Massey and Thaler,, 2013).
In this paper, we challenge that assertion. Massey and Thaler’s conclusion critically depends on a single questionable assumption, namely that general managers should maximize expected surplus value. We argue that this assumption is overly simplistic and does not necessarily capture the strategic objectives of NFL general managers. Winning in the NFL may depend less on accumulating surplus value and more on acquiring transformational players––those capable of fundamentally altering a team’s trajectory. If general managers instead prioritize maximizing the probability of drafting elite players, their decisions—–while appearing suboptimal under Massey and Thaler’s framework––may in fact be entirely rational.
In other words, we question whether expected value is the proper “rational” criterion for evaluating risky prospects. It is well known that individuals play the lottery and prior research has shown that individuals sometimes have a preference for positively skewed payoffs (e.g. Barberis and Huang, (2008)). However, a preference for positively skewed payoffs is often viewed as irrational in the context of traditional economic theories, which assume that individuals make decisions based on maximizing expected utility or wealth. The present research, in contrast, highlights an important decision context where, under reasonable assumptions, a preference for positively skewed payoffs may be perfectly rational.
This perspective provides a more nuanced view of decision making in high-stakes environments. Perhaps irrationality and cognitive biases are not as pervasive as Massey and Thaler, along with the broader judgement and decision making literature, assert. Our findings underscore the importance of specifying a decision-maker’s utility function before labeling behavior as irrational. More broadly, we highlight the risks of applying oversimplified economic models to complex, domain-specific decision-making contexts without considering alternative utility structures.
Below, we begin with an exposition of the setting of our case study––the NFL draft––and then give a more detailed overview of Massey and Thaler’s analysis and our rebuttal.
1.1 Overview
Players entering the NFL are allocated to teams through an annual draft. The draft consists of seven rounds, in which teams take turns selecting players in reverse order of the previous season’s win-loss records. The team with the worst record picks first, while the reigning Super Bowl champion selects last (with some minor exceptions). Trading draft picks, involving the exchange of current and future draft selections as well as active players, is a fundamental component of constructing an NFL roster. Teams typically engage in these trades to secure specific players or improve positional depth. For example, in 2013 the Miami Dolphins traded the $3^{rd}$ overall pick to the Oakland Raiders in exchange for the $12^{th}$ and $42^{nd}$ overall picks.111 https://bleacherreport.com/articles/1618345-miami-dolphins-select-dion-jordan-with-no-3-pick-in-trade-with-oakland-raiders Trades like this raise fundamental questions about the valuation of draft picks. How did the Dolphins and Raiders determine that their trade was mutually beneficial? More broadly, if a team seeks to acquire a specific draft pick, what constitutes an appropriate offer? If presented with a trade proposal involving multiple draft picks, how should a team evaluate its value and decide whether to accept or reject the exchange? A rigorous analysis of draft pick valuation is essential to understanding the trade dynamics that shape NFL roster management.
For decades, analysts and researchers have used draft position value curves to quantify the value of draft positions and assess the fairness of trades involving draft picks. Draft curves assign a numerical point value to each draft position $1,...,256$ and are typically additive: the value of a bundle of picks is the sum of the values of each individual pick. A trade is considered fair if the sum of the values on both sides is equal. In the 1990s, at the request of Dallas Cowboys head coach Jimmy Johnson, vice president Mike McCoy developed the first draft value curve222 https://www.the33rdteam.com/jimmy-johnsons-draft-trade-value-chart-set-standard-but-clearly-outdated/ based on intuition and an analysis of past trades. This curve assigns the highest value to the first overall pick and follows a steep, convex decline. For instance, in the 2013 draft trade between the Dolphins and the Raiders, Johnson’s curve suggests the Dolphins’ draft capital increased by 31%. The Raiders traded a single pick worth 2,200 points in exchange for two picks collectively worth 1,680 points (1,200 and 480, respectively). Although Johnson’s chart is ad hoc and lacks formal rigor, it remained widely used by general managers for decades, and some teams continue to rely on it today.
More recently, in their seminal work, Massey and Thaler, (2013) developed the first rigorous data-driven draft value curves, which are visualized in Figure 1. The solid line in Figure 1a denotes expected performance value, which quantifies the expected on-field performance of players drafted at a given position, expressed on a salary-equivalent scale. The dot-dashed line reflects first contract compensation, which denotes the cost of drafted players during their first contract. A player’s first contract compensation is essentially deterministic, as it is dictated by draft position. [^3] Drafted rookies may negotiate certain contract elements, such as performance bonuses or injury guarantees. However, this compensation is often much lower than the value of the player’s on-field performance, leading to what is known as first contract surplus value. The dashed lines in Figures 1a and 1b reflect expected surplus value, which measures the expected difference between a player’s first contract performance and his cost. Finally, the solid line in Figure 1b is the trade market curve, a draft value curve that best fits the dataset of all observed trades involving only draft picks. This curve reflects the draft value implied by actual trade market behavior.
Figure 1: Figure 3 from Massey and Thaler, (2013), which summarizes their results. In (a) they visualize expected performance value (solid line), first contract compensation (dot-dashed line), and expected surplus value (dashed line) versus draft position, expressed on a salary-equivalent scale. In (b) they visualize expected surplus value (dashed line) and the trade market curve (solid line) versus draft position, expressed relative to the value of the first pick.
Massey and Thaler identified two key insights. First, they discovered the loser’s curse, visualized in Figure 1a. Teams with the worst win-loss records in the previous season—the losers—receive the top draft picks, as expected performance value declines monotonically and convexly throughout the draft. However, these teams are cursed because top draft picks generate less surplus value on average than mid-to-late first-round picks, which typically belong to better teams. Expected surplus value (i.e., expected performance above cost) increases throughout the early first round, peaking in the middle of the round. Second, Massey and Thaler demonstrated that the draft pick trade market is highly inefficient, as shown in Figure 1b. The trade market curve is much steeper than the expected surplus value curve and differs in shape. Teams consistently overpay to trade up in the draft, investing excessive draft capital to select “top draft picks [that] are significantly overvalued” (Massey and Thaler,, 2013). In contrast, trading down and accumulating additional draft picks is an effective strategy to maximize surplus value. For example, consider the 2013 draft trade between the Dolphins and the Raiders. Using expected surplus value,444 This calculation is based on our expected surplus value curve, visualized in Figure 2 in Section 3, as the exact values from Massey and Thaler’s curve are unavailable. the Raiders’ draft capital increased by 135%. The Dolphins traded two picks collectively worth 2.40 points (1.51 and 0.89, respectively) in exchange for a single pick worth 1.02. Notably, the $12^{th}$ overall pick had a higher expected surplus value than the $3^{rd}$ overall pick (i.e., $1.51>1.02$). Although the Dolphins won the trade according to Jimmy Johnson’s draft value chart, the Raiders accumulated substantially more expected surplus value.
Massey and Thaler’s findings, summarized in Figure 1, are corroborated by recent data555 That Massey and Thaler’s results replicate to recent data was originally noticed as part the development of a forthcoming paper Avery et al., ming (see Figure 2 in Section 3). Despite the widespread recognition of their work over an extended period, the draft strategies employed by general managers in the NFL have remained largely unchanged. This persistence raises an important question: why has there been no apparent adjustment in decision making? One possible explanation––examined in the forthcoming paper Avery et al., ming––is that NFL general managers are resistant to change and continue to engage in suboptimal trade practices due to cognitive biases or misjudgment. However, this perspective prompts further inquiry: are these executives, responsible for managing multi-billion-dollar franchises, fundamentally misguided, or is there a crucial element missing from existing analyses? This paper explores the latter hypothesis, proposing an alternative explanation. Specifically, we argue that there may be a sound economic rationale underlying observed trade behavior. It is plausible that general managers are employing a different, yet entirely rational, utility function that does not rely solely on the expected future value of a player at a given draft position.
Perhaps the overarching strategic objective of NFL teams is to win the Super Bowl––a goal consistently emphasized by team owners, general managers, coaches, and players. As Kansas City Chiefs quarterback Patrick Mahomes succinctly stated, “the only goal is to win the Super Bowl.” [^6] https://x.com/_MLFootball/status/1826003252460269907?t=JvStTWyM92jEe21LxqOTPw&s=19 Massey and Thaler’s conclusions are predicated on the assumption that general managers, in pursuit of this objective, should adopt expected value as their guiding utility function. However, while their analysis relies on this assumption, they do not explicitly acknowledge the primacy of winning the Super Bowl or establish a direct link between maximizing expected surplus value and achieving that goal. While it is possible that maximizing expected surplus value contributes to winning, alternative utility functions may yield entirely different strategic conclusions. This paper develops that possibility, exploring the idea that Super Bowl success hinges on acquiring transformational players––those capable of profoundly altering a team’s trajectory––rather than optimizing expected surplus value. One example is the selection of Patrick Mahomes in the 2017 NFL Draft, which fundamentally elevated the Chiefs from a playoff-caliber team to a perennial Super Bowl contender.777 From 2013 to 2017, under coach Andy Reid but without starting quarterback Patrick Mahomes, the Chiefs made the playoffs in four of five seasons but won just one playoff game. From 2018 to 2023, the Chiefs had four Super Bowl appearances and won three. Another is Josh Allen, selected in 2018, who elevated the Buffalo Bills from years of ineptitude to one of the league’s top teams.888 From 2000 to 2017, the Bills made the playoffs once (in 2017). From 2018 to 2024, the Bills made the playoffs six times and made the AFC champsionship game twice. Under a different, but perhaps entirely reasonable, utility function––maximizing the probability of drafting transformational or elite players––decisions that appear suboptimal under Massey and Thaler’s framework may actually be entirely rational.
It is possible that the most effective strategy for constructing a Super Bowl caliber roster is not maximizing expected performance (or surplus) value but rather valuing draft positions based on the likelihood of acquiring elite talent. These two objective functions––maximizing expected performance (or surplus) value versus maximizing the expected number of elite players––differ significantly in their strategic implications. While both are based on expected values, the latter considers the probability that a player will produce exceptional, high-impact performance, thereby capturing the importance of tail outcomes in shaping team success.
A key insight, overlooked by previous researchers, is how variance in performance value changes over the course of the draft. Just as expected value declines convexly throughout the draft (see Figure 1a), so does variance (see Figure 3 in Section 4). We hypothesize that general managers are not only trading for expected value but also for variance, which is directly related to the probability of acquiring a transformational player in the far right tail of performance. For example, consider the 2013 draft trade between the Dolphins and the Raiders. In this exchange, the Dolphins traded expected value to the Raiders in exchange for variance. Based on the expected number of elite players,999 Here, we define an elite player as one at least as valuable as Jared Goff, represented by the orange curve in Figure 14 in Section 6. the Dolphins’ draft capital increased by 40%. The Raiders traded a single pick worth 0.893 expected elite players for two picks collectively worth 0.639 (0.538 and 0.101). Although the Raiders accumulated substantially more expected surplus value, the Dolphins increased their expected number of elite players. Further, notice how much less the $42^{nd}$ pick is worth than before––it produces 89% fewer expected elite players than the $3^{rd}$ pick, but just 13% less expected surplus value. The likelihood of acquiring transformational players decays precipitously across the draft.
Hence, in this paper we construct Bayesian hierarchical draft value curves that reflect the probability of drafting an elite player. We find that general managers’ behavior in the trade market aligns with a decision rule that prioritizes variance and right-tail outcomes––seeking players with the potential to be game-changing superstars, even if they are riskier investments. Quarterbacks in particular follow a more extreme version of this pattern. Under the lens of drafting elite players, particularly for teams seeking a franchise quarterback, general managers’ behavior in the draft pick trade market appears rational rather than systematically flawed.
The remainder of this paper is organized as follows. We begin in Section 2 by defining the value of an NFL player’s performance. Then, in Section 3, we review the traditional approach to constructing draft position value curves, similar to those devised in Massey and Thaler, (2013). In Section 4 we elaborate on our research hypothesis, describing how general managers’ decision making may be rational under alternative utility functions that prioritize variance and right-tail outcomes over expected surplus value. To construct other draft position value curves built from alternative value functions, we estimate the full conditional density of player performance given draft position and player position in Section 5. We present our results in Section 6 and conclude in Section 7.
2 Measuring player performance
@@
Since the value of a draft pick is realized by the on-field performance of the player it produces, a crucial first step is to measure the value of a player’s performance. Doing so takes care and consideration. Each play in American football consists of a complex interplay between 22 players of varying roles interacting in continuous time and space. This makes it difficult to measure how a player’s individual performance impacts the ultimate outcome of a play. For instance, a linebacker may impact a play simply by occupying space. He could cause the ball carrier to move in a different direction, and there is no easy way to measure the impact of that action. There is no industry-standard way of measuring player performance on a given play in American football, particularly in a way that measures players of all positions on the same scale. Hence, analysts use performance proxies. Each measure is imperfect and has its own drawbacks. We describe many of these performance measures in detail in Appendix B. Throughout this paper, for simplicity, we use one performance outcome variable, detailed below.
@@
In free agency, NFL players who are not bound by a contract can sign a new contract with any team. We believe the free agent market is relatively efficient and thus reflects the performance value of players. Therefore, following Baldwin, (2024), in this study we measure performance value by the average salary per year of a contract signed in free agency. We consider salary as a percentage of the cap––a rule that determines the maximum amount any team can pay its players in a given year––to adjust for inflation. Over time the cap has increased from \$123 million in 2013 to \$224.8 million in 2023.
@@
Researchers, beginning with Massey and Thaler, (2013), have been interested in particular in the value of a player’s performance during his first contract. This is because an NFL player is paid particularly cheaply during his first contract. A player’s first contract is determined by his draft position and can be much lower than the value of his performance, creating a first contract “surplus value.” Teams can take advantage of the resulting extra salary cap space by signing good veteran players in free agency. The drawback of focusing solely on a player’s first contract is that the very best players stay with the same team across multiple contracts, providing value across their entire careers. For example, Tom Brady played in the NFL for 23 seasons and won seven Super Bowls. Continuing with the precedent set by Massey and Thaler, (2013) and subsequent analyses, in this study we consider just first contract performance and surplus value.
@@
Throughout this paper, we measure a player’s first contract performance by his second contract’s average salary per year as a percentage of the cap. This is a reflection of the value of a player’s first contract performance because he signs his second contract in free agency just after his first contract ends. Players who don’t sign a second contract have zero performance value. We acknolwedge that free agent annualized cap percentage is an imperfect measure of a player’s performance (see Appendix B) and first contract value is an incomplete measure of the total value he provides to his team (as it excludes the rest of his career). Nonetheless, the focus of this paper is to value draft picks given a performance measure, not to choose the best performance measure. We leave further analysis of various measures of performance to future work.
3 Traditional draft position value curves
@@
Now, we review the traditional approach to constructing draft position value curves. Prior to the draft, we do not know how well a player drafted at draft position $x\in\{1,...,256\}$ will perform in the NFL. Thus, we think of the performance outcome $Y$ associated with position $x$ as a random variable, denoted by a capital letter. Recall that, throughout this paper, we let the realization of $Y$ be a player’s annualized second contract value as a percentage of the salary cap. Then, from the dataset $\{(x_{i},Y_{i})\}$ of all recent draft picks in NFL history, we want to learn the de-noised relationship between pick number $x$ and performance outcome $Y$. Traditional analyses use a standard approach but with different measures of performance, detailed in Appendix B (Massey and Thaler,, 2013; Stuart,, 2012; Fitzgerald and Spielberger,, 2024; Pro Football Focus,, 2024; Baldwin,, 2024). The de-noised relationship they estimate from data is the expected value of $Y$ given $x$.
@@
This traditional approach proceeds as follows. First, fit the expected performance value curve, or the conditional mean $x\mapsto{\mathbb{E}}[Y|x]$. Second, calculate the cost curve, $x\mapsto{\mathrm{cost}}(x)$. Cost is first contract compensation, which is essentially a deterministic function of $x$. Third, calculate the expected surplus value curve, or the expected value of the difference between performance and cost, $x\mapsto{\mathbb{E}}[Y-{\mathrm{cost}}(x)|x]$. Finally, normalize each of these curves by dividing by the value of the first pick. Value is then relative to the first pick, whose value is one.
@@
We fit these traditional draft value curves using our outcome variable $Y$ (second contract annualized cap percentage), which we visualize in Figure 2. We use spline regression to fit the expected performance value curve from a dataset of all draft picks from 2013 to 2023 (Ho and Carl,, 2024). We calculate the cost curve as in Baldwin, (2024): beginning with the actual 2023 salary cap (about \$$\$225$ million) and assuming a cap growth rate of 7% over the next three seasons (2024, 2025, and 2026), we convert the total dollar amount of the rookie contract into a percentage of the cap. Then, for each draft position we average the compensation values across the four seasons. We calculate the expected surplus value curve as the difference between the expected performance value curve and cost curve. Finally, we normalize each of these curves so that value is relative to the first pick. Because of this normalization, note that the expected surplus value curve in Figure 2 does not visually look like the expected performance value curve minus the cost curve. In Figure 2 we also visualize a draft value curve implied by the trade market. Devised in Massey and Thaler, (2013), the fitted trade market curve is a singular draft value curve that best fits the dataset of all observed trades. We derive this curve in detail in Appendix C.
@@
Figure 2: Each curve is a valuation of draft position $x$ relative to the first pick. The dotted black line is the expected performance value curve, proportional to $x\mapsto{\mathbb{E}}[Y|x]$. The solid gray line is the cost curve, proportional to $x\mapsto{\mathrm{cost}}(x)$. The long-dashed black line is the expected surplus value curve, proportional to $x\mapsto{\mathbb{E}}[Y-{\mathrm{cost}}(x)|x]$. The solid black line is the draft value curve implied by the trade market.
@@
Expected performance value is less steep than cost, indicating that players are underpaid on average in their first contracts. This produces positive expected surplus value across most of the draft (the long-dashed black line is mostly positive). Expected surplus value is larger than one throughout the first round, peaking in the middle of the first round. This surplus spike is the “loser’s curse” from Massey and Thaler, (2013): top draft picks, belonging to last year’s worst teams, produce less surplus value on average than draft picks later in the first round that belong to better teams.